# Left-left-right-right magnetic order in spin-1/2 Kitaev-Heisenberg chain

### v1

@@ IV LLRR phase in the Kitaev-Heisenberg-Gamma chain
In this section, we apply the perturbative Luttinger liquid analysis to the Kitaev-Heisenberg-Gamma chain close to the FM point, by adding a nonzero Gamma term. Because of the equivalence in Eq. (6), we will focus on the $\Gamma>0$ region.
@@
where $D^{(\text{I})}_{2}$ is still given by Eq. (45). The most general magnetic pattern with an unbroken symmetry group $H^{(\text{I})}_{1}$ in the $U_{4}$ frame has been determined in Ref. Yang2020, as
@@
This is a distorted spiral order, which gains components in the $z$-direction compared with the magnetic ordering in Eq. (48) for $\Gamma=0$. Clearly, Eq. (48) can be obtained from Eq. (62) by setting $f=k=\frac{1}{\sqrt{2}}N_{1}$ and $h=0$.
@@
where $D^{(\text{II})}_{2}$ is still given by Eq. (45). The most general magnetic pattern with an unbroken symmetry group $H^{(\text{II})}_{1}$ in the $U_{4}$ frame can be determined as
$\vec{S}^{\prime(\text{II})}_{1}$ $=$ $(0,b,-c)^{T},$
$\vec{S}^{\prime(\text{II})}_{2}$ $=$ $(0,b,-c)^{T},$
$\vec{S}^{\prime(\text{II})}_{3}$ $=$ $(0,b,c)^{T},$
$\vec{S}^{\prime(\text{II})}_{4}$ $=$ $(0,b,c)^{T}.$ (65)
@@
$\vec{S}_{1}^{(\text{II})}$ $=$ $(0,b,c)^{T},$
$\vec{S}_{2}^{(\text{II})}$ $=$ $(0,-b,-c)^{T},$
$\vec{S}_{3}^{(\text{II})}$ $=$ $(0,-b,-c)^{T},$
$\vec{S}_{4}^{(\text{II})}$ $=$ $(0,b,c)^{T}.$ (66)
This is a distorted left-left-right-right magnetic order when $a$ and $b$ have the same signs having components along the $z$-directino as well. Clearly, Eq. (57) can be obtained from Eq. (66) by setting $b=\frac{1}{\sqrt{2}}N_{2}$, $c=0$.
Finally, we note that the $c=1$ phase in the Kitaev-Heisenberg model extends to a Kitaev spin liquid phase when $\Gamma$ is nonzero as shown in Fig. 1 (c) (denoted as “KSL” in the figure), whose nature remains unclear and will be left for future investigations.
V The FM phase in the $J<0$ region
Although this work mainly focuses on the LLRR phase in the $J>0$ region, in this section we briefly discuss the FM phase in the $J<0$ region in Fig. 1 in the framework of perturbative Luttinger liquid analysis.
@@ V The FM phase in the $J<0$ region
For the Kitaev-Heisenberg model, the region between $-K$ and $-J$ in Fig. 1 (a) is unitarily equivalent to the region between $-K$ and FM under the $U_{4}$ transformation. Hence, in the neighborhood of $-J$, the system is in an ordered phase, and the magnetic pattern in the original frame in this case is exactly the same as the pattern in the $U_{4}$ frame in the vicinity of FM. Therefore, in the region $[\phi_{c2},\pi]$ in Fig. 1 (a) where $\phi_{c2}$ is dual to $\phi_{c1}$ under the equivalent relation in Eq. (4), we conclude that for $g>0$, the magnetic ordering is along $\frac{1}{\sqrt{2}}(\pm 1,\pm 1,0)$-directions (see Eq. (47) and the other three degenerate solutions); and for $g<0$, the magnetic ordering is along ±$\pm\hat{x}$- and ±$\pm\hat{y}$-directions (see Eq. (56) and the other three degenerate solutions).
@@
$\tilde{H}_{XXZ}=|J+\frac{1}{2}K|\sum_{i}[\tilde{S}_{i}^{x}\tilde{S}_{i+1}^{x}+\tilde{S}_{i}^{y}\tilde{S}_{i+1}^{y}+\Delta\tilde{S}_{i}^{z}\tilde{S}_{i+1}^{z}],$ (68)
$H_{K}=\frac{1}{2}K\sum_{i}(-)^{i}(\tilde{S}_{i}^{x}\tilde{S}_{i+1}^{x}-\tilde{S}_{i}^{y}\tilde{S}_{i+1}^{y}),$ (69)
@@
in which $\tilde{S}_{i}$ represents the spin operators after the $V_{2}$ transformation, and
@@
$H^{(\text{I})}=\langle T_{a}I,R(\hat{n}_{1},\pi)T_{a}\rangle,$ (78)
and the most general from of the magnetic pattern invariant under $H^{(\text{I})}$ can be determined as
@@
Notice that the magnetic ordering in Eq. (82) only represents a special configuration consistent with the unbroken symmetry group $H^{(\text{II})}$. In fact, the most general from of the magnetic pattern invariant under $H^{(\text{II})}$ can be determined as
@@
VI DMRG numerics
In this section, we provide DMRG numerical evidences for the predictions in the previous sections.
We start with the Kitaev-Heisenberg model discussed in Sec. III. Fig. 5 shows the values of the central charge in the region $\phi\in(\arctan(2),\pi/2)$ (where $J=\cos(\phi)$, $K=-\sin(\phi)$) extracted from entanglement entropy Calabrese2009 at different system sizes with open boundary conditions. It can observed from Fig. 5 that when $\phi<0.4\pi$, the central charge value decreases by increasing system size, having a tendency to approach zero in the limit of infinite sizes. This is consistent with the analytic prediction in Sec. III on the existence of a magnetically ordered phase when $\phi$ is close to $\arctan(2)$. On the other hand, when $\phi>0.4\pi,$the value of the numerically extracted central charge increases by increasing the system size, tending to approach the value of $1$ in the large size limit, thereby indicating a $c=1$ gapless phase in the region $\phi\in(\arctan(2),\pi/2)$. The low energy field theory describing this gapless phase is non-perturbative in nature, which remains unclear and will be left for future studies.
Figure 4: Central charge values of the spin-1/2 Kitaev-Heisenberg chain as a function of $\phi$. DMRG numerics are performed for $L=48,100,256$ sites using open boundary conditions.
Notice that there are two possible types of magnetic orders according to the analysis in Sec. III and Sec. IV, namely, the spiral order and the LLRR order, both in the unrotated frame. In order to determine which type of magnetic order develops in the system, we study the response of the spin expectation values to a small uniform external magnetic field. When there is spontaneous symmetry breaking, it is known that an infinitesimal field is enough to produce finite spin expectation values in the thermodynamic limit.
For convenience, we perform DMRG numerics in the four-sublattice rotated frame, and accordingly, the patterns of spin expectation values of the spiral phase and the LLRR phase in the $U_{4}$ frame are given by Eq. (61) and Eq. (65), respectively. When a uniform $(1,1,0)$-field is applied in the $U_{4}$ frame, it is expected that the spin pattern in Eq. (61) is developed. Fig. 5 displays the numerical results of the spin expectation values at $(\theta=0.38\pi,\phi=0.98\pi)$ when a $h=2\times 10^{-4}$ field is applied along the $(1,1,0)$-direction. As can be seen from Fig. 5, the pattern is consistent with Eq. (61). On the other hand, when a uniform $(0,1,0)$-field is applied in the $U_{4}$ frame, it is expected that the pattern in Eq. (65) develops in the system. Fig. 6 displays the numerical results of the spin expectation values at $(\theta=0.38\pi,\phi=0.98\pi)$ when a $h=2\times 10^{-4}$ field is applied along the $(0,1,0)$-direction. It can be observed from Fig. 6 that the pattern is consistent with Eq. (65).
Figure 5: Spin expectation values $\langle S^{x}\rangle$ (marked with blue star), $\langle S^{y}\rangle$ (marked with red circle), and $\langle S^{z}\rangle$ (marked with yellow square) as functions of site $i$ when a magnetic field $h=2\times 10^{-4}$ is applied along the $(1,1,0)$-direction. DMRG numerics are carried out on a system of $L=96$ sites using periodic boundary conditions at $(\theta=0.38\pi,\phi=0.98\pi)$ with a truncation error $\epsilon=10^{-10}$.
Figure 6: Spin expectation values $\langle S^{x}\rangle$ (marked with blue star), $\langle S^{y}\rangle$ (marked with red circle), and $\langle S^{z}\rangle$ (marked with yellow square) as functions of site $i$ when a magnetic field $h=2\times 10^{-4}$ is applied along the $y$-direction. DMRG numerics are carried out on a system of $L=96$ sites using periodic boundary conditions at $(\theta=0.38\pi,\phi=0.98\pi)$ with a truncation error $\epsilon=10^{-10}$. The data for $\langle S^{x}\rangle$ and $\langle S^{y}\rangle$ overlap and cannot be distinguished.
To determine which type of magnetic orders wins, we study the change $\Delta E$ in the ground state energy when magnetic fields in different directions are applied to the system. Fig. 7 shows the numerical results of the energy response to applied fields in the $(1,1,0)$- (blue) and $(0,1,0)$- (red) directions at $(\theta=0.38\pi,\phi=0.98\pi)$. As can be seen from Fig. 7, the system has a larger response to the $(0,1,0)$-field. Hence, it is the LLRR order that is developed in the system in the unrotated frame, rather than the spiral order.
Figure 7: Energy change $\Delta E$ as a function of magnetic field along $(1,1,0)$- (blue) and $(0,1,0)$- (red) directions. DMRG numerics are performed on systems of $L=144$ sites using periodic boundary conditions at $(\theta=0.38\pi,\phi=0.98\pi)$.
VII Weakly coupled chains: 2D stripy order
@@
VIII Summary
In summary, we have studied the phase diagram of the spin-1/2 Kitaev-Heisenberg chain in the $(K<0,J>0)$ region, based on a combination of perturbative Luttinger liquid analysis and DMRG numerics. Close to the hidden SU(2) symmetric FM point, an ordered phase is identified having a left-left-right-right magnetic order, whereas close to the FM Kitaev point, a gapless phase with central charge value $c=1$ is found. The Luttinger liquid analysis is extended to the Kitaev-Heisenberg-Gamma model as well as the $(K<0,J<0)$ region. Furthermore, the 2D stripy phase is recovered based on the 1D left-left-right-right order using a coupled-chain method, thereby providing a quasi-1D explanation for the stripy order on the honeycomb lattice.
@@ Acknowledgements.
W.Y. is supported by the startup funding of Nankai University. C.X. acknowledges the supports from MOST grant No. 2022YFA1403900, NSFC No. 12104451, NSFC No. 11920101005, and funds from Strategic Priority Research Program of CAS No. XDB28000000. A.N. acknowledges computational resources and services provided by Compute Canada and Advanced Research Computing at the University of British Columbia. A.N. acknowledges support from the Max Planck-UBC-UTokyo Center for Quantum Materials and the Canada First Research Excellence Fund (CFREF) Quantum Materials and Future Technologies Program of the Stewart Blusson Quantum Matter Institute (SBQMI).

### v2

@@ IV Distorted-LLRR phase in Kitaev-Heisenberg-Gamma chain
In this section, we apply the perturbative Luttinger liquid analysis to the Kitaev-Heisenberg-Gamma chain close to the hidden SU(2) symmetric FM point, by adding a nonzero Gamma term. Because of the equivalence in Eq. (6), we will focus on the $\Gamma>0$ region.
IV.1 Perturbative Luttinger liquid analysis
@@
where $D^{(\text{I})}_{2}$ is still given by Eq. (45). Combining Eq. (14) with Eq. (67), it can be observed that the symmetry breaking pattern is
$D_{4h}\rightarrow D_{2}^{\text{(I)}}.$ (68)
The most general magnetic pattern with an unbroken symmetry group $H^{(\text{I})}_{1}$ in the $U_{4}$ frame has been determined in Ref. Yang2020, as (for derivations, see Appendix B.3)
@@
This is a distorted spiral order, which gains components in the $z$-direction compared with the magnetic ordering in Eq. (48) for $\Gamma=0$. Clearly, Eq. (48) can be obtained from Eq. (70) by setting $f=k=\frac{1}{\sqrt{2}}N_{1}$ and $h=0$. The pattern of the magnetic ordering for the other three degenerate ground states can be obtained by applying the broken symmetry transformations to Eq. (70).
@@
where $D^{(\text{II})}_{2}$ is still given by Eq. (45). Combining Eq. (14) with Eq. (72), it can be observed that the symmetry breaking pattern is
$D_{4h}\rightarrow D_{2}^{\text{(II)}}.$ (73)
The most general magnetic pattern with an unbroken symmetry group $H^{(\text{II})}_{1}$ in the $U_{4}$ frame can be determined as (for derivations, see Appendix B.4)
$\vec{S}^{\prime(\text{II})}_{1}$ $=$ $(a,b,-c)^{T},$
$\vec{S}^{\prime(\text{II})}_{2}$ $=$ $(-a,b,-c)^{T},$
$\vec{S}^{\prime(\text{II})}_{3}$ $=$ $(a,b,c)^{T},$
$\vec{S}^{\prime(\text{II})}_{4}$ $=$ $(-a,b,c)^{T}.$ (74)
@@
$\vec{S}_{1}^{(\text{II})}$ $=$ $(-a,b,c)^{T},$
$\vec{S}_{2}^{(\text{II})}$ $=$ $(a,-b,-c)^{T},$
$\vec{S}_{3}^{(\text{II})}$ $=$ $(a,-b,-c)^{T},$
$\vec{S}_{4}^{(\text{II})}$ $=$ $(-a,b,c)^{T}.$ (75)
This is a distorted left-left-right-right magnetic order. Clearly, Eq. (58) can be obtained from Eq. (75) by setting $b=\frac{1}{\sqrt{2}}N_{2}$, $c=0$. The pattern of the magnetic ordering for the other three degenerate ground states in this case can be obtained by applying the broken symmetry transformations to Eq. (75).
Finally we note that the distorted-LLRR phase in the Kitaev-Heisenberg-Gamma model and the LLRR phase in the Kitaev-Heisenberg model belong to the same phase, since the broken symmetries in both cases are the same, as can be inspected from Eq. (56) and Eq. (73). Similar conclusion holds for the distorted-spiral and spiral phases.
IV.2 DMRG numerics
In this subsection, we provide DMRG numerical evidence for distorted-LLRR order in the Kitaev-Heisenberg-Gamma chain by including a small Gamma term. The following parametrization will be used for $K,J,\Gamma$ in this subsection
$J$ $=$ $\cos(\theta)$
$K$ $=$ $\sin(\theta)\cos(\phi)$
$\Gamma$ $=$ $\sin(\theta)\sin(\phi),$ (76)
in which $\theta\in[0,\pi]$, $\phi\in[0,2\pi]$. Notice that in terms of Eq. (76), the point $(\theta,\pi)$ corresponds to the Kitaev-Heisenberg model at angle $\phi$ using the parametrization in Eq. (2). We will take a representative point $(\theta=0.37\pi,\phi=0.99\pi)$ in DMRG numerical calculations in this subsection, which has a small nonzero value of $\Gamma$ since $\phi$ is close to $\pi$. We stick to the four-sublattice rotated frame in all numerical calculations in this subsection.
IV.2.1 Magnetic ordering
Figure 10: Spin expectation values in the Kitaev-Heisenberg-Gamma model throughout the chain at $(\theta=0.37\pi,\phi=0.99\pi)$ in the presence of a small field of strength $2\times 10^{-4}$ along (a) $(1,1,0)$-direction, and (b) $(0,1,0)$-direction. DMRG numerics are performed on a system of $L=96$ sites with periodic boundary conditions. In (a), the data points for expectation values along $x$ and $y$ spin directions coincide; and the inset figure is a zoom-in view of the spin expectation values along $x$- and $y$-directions. The bond dimension $m$ and truncation error $\epsilon$ in DMRG calculations are taken as $m=1200$, $\epsilon=10^{-10}$.
In the Kitaev-Heisenberg-Gamma model, there are again two types of magnetic orders according to the analysis in Sec. IV.1, similar to the Kitaev-Heisenberg case. To test the magnetic order, we numerically calculate spin expectation values by applying small magnetic fields along the $(1,1,0)$- and $(0,1,0)$-directions.
DMRG numerics are performed in the $U_{4}$ frame at $(\theta=0.37\pi,\phi=0.99\pi)$ for small fields $h_{xy}=2\times 10^{-4}$ and $h_{y}=2\times 10^{-4}$ along $(1,1,0)$- and $(0,1,0)$-directions, respectively, and the results for spin expectation values are shown in Fig. 10. As can be observed, the patterns of spin expectation values in Fig. 10 (a) for $h_{xy}$ and Fig. 10 (b) for $h_{y}$ are consistent with the predictions in the $U_{4}$ frames for the distorted-spiral order in Eq. (69) and the distorted-LLRR order in Eq. (74), respectively. Similar to the case of Kitaev-Heisenberg model, there is the question which one is the actual order in the system, which will be answered shortly by the study of energy-field relation in Sec. IV.2.2.
IV.2.2 Energy-field relation
Fig. 11 shows DMRG numerical results for ground state energy per site $E_{0}/L$ at $(\theta=0.36\pi,\phi=0.99\pi)$ as a function of the applied small fields along $(1,1,0)$- and $(0,1,0)$-directions in the $U_{4}$ frame for two system sizes $L=48$ and $96$. It is clear that the relation between energy change and applied fields are linear in Fig. 11, confirming the existence of a magnetic order. As can be inspected in Fig. 11, the energy changes $\Delta E_{y}$ is indeed larger than $\Delta E_{xy}$ when $h_{y}=h_{xy}$, consistent with the distorted-LLRR order rather than the distorted-spiral order according to a similar analysis as Sec. III.4.4.
Figure 11: Ground state energy per site $E_{0}/L$ in the Kitaev-Heisenberg-Gamma model as a function of applied small fields $h_{xy}$ (dashed lines) and $h_{y}$ (dashed lines), where $h_{xy}$ and $h_{y}$ are along $(1,1,0)$- and $(0,1,0)$-directions, respectively. DMRG numerics are performed on system sizes $L=48$ (blue lines) and $L=96$ (red lines) at $(\theta=0.37\pi,\phi=0.99\pi)$ using periodic boundary conditions. The bond dimension $m$ and truncation error $\epsilon$ in DMRG calculations are taken as $m=1200$, $\epsilon=10^{-10}$.
V Other phases in Kitaev-Heisenberg-Gamma chain
In this section, we make comments on other phases in the Kitaev-Heisenberg-Gamma model in Fig. 1, especially the FM phase. The antiferromagnetic phase (denoted as “AFM”), the Luttinger liquid phase (denoted as “LL”), and the ferromagnetic phase (denoted as “FM”) have been investigated in details both analytically and numerically in Ref. Yang2020, . Interestingly, the FM phase in the $J<0$ region can also be derived in the framework of perturbative Luttinger liquid analysis in a similar manner as Sec. IV.1, which will be shown below.
V.1 The FM phase
@@
For the Kitaev-Heisenberg model, the region between $-K$ and $-J$ on the horizontal axis in Fig. 1 is unitarily equivalent to the region between $-K$ and FM under the $U_{4}$ transformation. Hence, in the neighborhood of $-J$, the system is in an ordered phase, and the magnetic pattern in the original frame in this case is exactly the same as the pattern in the $U_{4}$ frame in the vicinity of FM. Therefore, in the region $[\phi_{c2},\pi]$ on the horizontal axis in Fig. 1 where $\phi_{c2}$ is dual to $\phi_{c1}$ under the equivalent relation in Eq. (4), we conclude that for $g>0$, the magnetic ordering is along $\frac{1}{\sqrt{2}}(\pm 1,\pm 1,0)$-directions (see Eq. (47) and the other three degenerate solutions); and for $g<0$, the magnetic ordering is along ±$\pm\hat{x}$- and ±$\pm\hat{y}$-directions (see Eq. (57) and the other three degenerate solutions).
@@
$\tilde{H}_{XXZ}$ $=$ $|J+\frac{1}{2}K|\sum_{i}[\tilde{S}_{i}^{x}\tilde{S}_{i+1}^{x}+\tilde{S}_{i}^{y}\tilde{S}_{i+1}^{y}+\Delta\tilde{S}_{i}^{z}\tilde{S}_{i+1}^{z}]$
$H_{K}$ $=$ $\frac{1}{2}K\sum_{i}(-)^{i}(\tilde{S}_{i}^{x}\tilde{S}_{i+1}^{x}-\tilde{S}_{i}^{y}\tilde{S}_{i+1}^{y})$
@@
where $\tilde{S}_{i}$ represents the spin operators after the $V_{2}$ transformation, and
@@
$H^{(\text{I})}_{2}=\langle T_{a}I,R(\hat{n}_{1},\pi)T_{a}\rangle,$ (86)
and the most general from of the magnetic pattern invariant under $H^{(\text{I})}$ can be determined as (for derivations, see Appendix B.5)
@@
Notice that the magnetic ordering in Eq. (90) only represents a special configuration consistent with the unbroken symmetry group $H^{(\text{II})}$. In fact, the most general from of the magnetic pattern invariant under $H^{(\text{II})}$ can be determined as (for derivations, see Appendix B.6)
@@
V.2 The KSL phase
We note that it remains unclear whether the $c=1$ phase in the Kitaev-Heisenberg model extends to a Kitaev spin liquid phase when $\Gamma$ is nonzero in Kitaev-Heisenberg-Gamma model as shown in Fig. 1 (denoted as “KSL” in the figure, which is “Kitaev spin liquid for short”), which will be left for future investigations.
VI Weakly coupled chains: 2D stripy order
@@
VII Summary
In summary, we have studied the phase diagram of the spin-1/2 Kitaev-Heisenberg chain in the $(K<0,J>0)$ region, based on a combination of perturbative Luttinger liquid analysis and DMRG numerics. Close to the hidden SU(2) symmetric FM point, an ordered phase is identified having a left-left-right-right magnetic order, whereas close to the FM Kitaev point, a gapless phase with central charge value $c=1$ is found. The Luttinger liquid analysis is extended to the Kitaev-Heisenberg-Gamma model as well as the parameter region of $(K<0,J<0)$. Furthermore, the 2D stripy phase is recovered based on the 1D left-left-right-right order using a coupled-chain method, thereby providing a quasi-1D explanation for the stripy order on the honeycomb lattice.
@@ Acknowledgements.
W.Y. is supported by the National Natural Science Foundation of China (Grants No. 12474476). C.X. acknowledges the supports from MOST grant No. 2022YFA1403900, NSFC No. 12104451, NSFC No. 11920101005, and funds from Strategic Priority Research Program of CAS No. XDB28000000. A.N. acknowledges computational resources and services provided by Compute Canada and Advanced Research Computing at the University of British Columbia. A.N. acknowledges support from the Max Planck-UBC-UTokyo Center for Quantum Materials and the Canada First Research Excellence Fund (CFREF) Quantum Materials and Future Technologies Program of the Stewart Blusson Quantum Matter Institute (SBQMI). The DMRG calculations in this work were performed using the software package ITensor in Ref. ITensor, .
@@ Appendix A Symmetry transformation properties of the bosonization fields
Appendix B Determinations of magnetic ordering patterns
In this appendix, we determine the patterns of magnetic orders using the unbroken symmetry groups, by requiring that the spin expectation values be invariant under the unbroken symmetries.
B.1 Spiral order in Kitaev-Heisenberg chain
The unbroken symmetry group $H^{\text{(I)}}$ for the spiral order in Kitaev-Heisenberg chain in the $U_{4}$ frame is given in Eq. (42), containing four symmetry elements. By considering a four-site unit cell, it is enough to consider the four generators $T_{a}I$, $T_{2a}$, $[R(\hat{z},-\frac{\pi}{2})T_{a}]^{2}T$, and $R(\hat{z},-\frac{\pi}{2})T_{a}\cdot R(\hat{y},\pi)T_{a}I$ (all modulo $T_{4a}$) for the group $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\ltimes D^{(\text{I})}_{2}$ in Eq. (43). It can be verified that under the four generators, the transformations of the spin expectation values $\langle\vec{S}_{i}\rangle=(x_{i},y_{i},z_{i})^{T}$ are
$T_{a}I$ ::: $\left(\begin{array}[]{c}x_{i}\\ y_{i}\\ z_{i}\end{array}\right)\rightarrow\left(\begin{array}[]{c}x_{1-i}\\ y_{1-i}\\ z_{1-i}\end{array}\right),$ (138)
$T_{2a}$ ::: $\left(\begin{array}[]{c}x_{i}\\ y_{i}\\ z_{i}\end{array}\right)\rightarrow\left(\begin{array}[]{c}x_{i+2}\\ y_{i+2}\\ z_{i+2}\end{array}\right),$ (145)
$(R(\hat{z},-\frac{\pi}{2})T_{a})^{2}T$ ::: $\left(\begin{array}[]{c}x_{i}\\ y_{i}\\ z_{i}\end{array}\right)\rightarrow\left(\begin{array}[]{c}x_{i+2}\\ y_{i+2}\\ -z_{i+2}\end{array}\right),$ (152)
$R(\hat{z},-\frac{\pi}{2})T_{a}R(\hat{y},\pi)T_{a}I$ ::: $\left(\begin{array}[]{c}x_{i}\\ y_{i}\\ z_{i}\end{array}\right)\rightarrow\left(\begin{array}[]{c}y_{2-i}\\ x_{2-i}\\ -z_{2-i}\end{array}\right).$ (159)
Invariance under $(R(\hat{z},-\frac{\pi}{2})T_{a})^{2}T$ requires
$x_{1}=x_{3},\leavevmode\nobreak\ y_{1}=y_{3},\leavevmode\nobreak\ z_{1}=-z_{3}$
$x_{2}=x_{4},\leavevmode\nobreak\ y_{2}=y_{4},\leavevmode\nobreak\ z_{2}=-z_{4}.$ (160)
Invariance under $R(\hat{z},-\frac{\pi}{2})T_{a}R(\hat{y},\pi)T_{a}I$ requires
$x_{1}=y_{1},\leavevmode\nobreak\ z_{1}=0$
$x_{3}=y_{3},\leavevmode\nobreak\ z_{3}=0$
$x_{2}=y_{4},\leavevmode\nobreak\ y_{2}=x_{4},\leavevmode\nobreak\ z_{2}=-z_{4}.$ (161)
Combined with invariance under $T_{a}I$ and $T_{2a}$, we obtain
$z_{1}=z_{2}=z_{3}=z_{4}=0$
$x_{1}=x_{2}=x_{3}=x_{4}=y_{1}=y_{2}=y_{3}=y_{4}.$ (162)
B.2 LLRR order in Kitaev-Heisenberg chain
The unbroken symmetry group $H^{\text{(II)}}$ for the spiral LLRR in Kitaev-Heisenberg chain in the $U_{4}$ frame is given in Eq. (52), generated by four elements $T_{a}I$, $T_{2a}$, $[R(\hat{z},-\frac{\pi}{2})T_{a}]^{2}T$, and $R(\hat{y},\pi)T_{a}I$. The actions of $T_{a}I$, $T_{2a}$, and $[R(\hat{z},-\frac{\pi}{2})T_{a}]^{2}T$ on spin expectation values have been given in Eq. (159), and the action of $R(\hat{y},\pi)T_{a}I$ is
$R(\hat{y},\pi)T_{a}I$ ::: $\left(\begin{array}[]{c}x_{i}\\ y_{i}\\ z_{i}\end{array}\right)\rightarrow\left(\begin{array}[]{c}-x_{1-i}\\ y_{1-i}\\ -z_{1-i}\end{array}\right).$ (169)
Invariance under $R(\hat{y},\pi)T_{a}I$ requires
$x_{1}=-x_{4},\leavevmode\nobreak\ y_{1}=y_{4},\leavevmode\nobreak\ z_{1}=-z_{4}$
$x_{2}=-x_{3},\leavevmode\nobreak\ y_{2}=y_{3},\leavevmode\nobreak\ z_{2}=-z_{3}.$ (170)
Combining the four generators, the magnetic ordering invariant under $H^{\text{(II)}}$ is given by
$x_{1}=x_{2}=x_{3}=x_{4}=0$
$y_{1}=y_{2}=y_{3}=y_{4}$
$z_{1}=z_{2}=z_{3}=z_{4}=0.$ (171)
This is the pattern given in Eq. (57).
B.3 Distorted spiral order in Kitaev-Heisenberg-Gamma chain
In the Kitaev-Heisenberg-Gamma chain, the unbroken symmetry group $H_{1}^{\text{(I)}}$ for the distorted spiral order is generated by $[R(\hat{z},-\frac{\pi}{2})T_{a}]^{2}T$ and $R(\hat{z},-\frac{\pi}{2})T_{a}\cdot R(\hat{y},\pi)T_{a}I$, hence only invariances in Eq. (160,160) are required, leading to
$x_{1}=y_{1}=x_{3}=y_{3},\leavevmode\nobreak\ z_{1}=z_{3}=0$
$x_{2}=y_{2}=x_{4}=y_{4},\leavevmode\nobreak\ z_{2}=-z_{4}.$ (172)
This is the pattern given in Eq. (69).
B.4 Distorted LLRR order in Kitaev-Heisenberg-Gamma chain
In the Kitaev-Heisenberg-Gamma chain, the unbroken symmetry group $H_{1}^{\text{(II)}}$ for the distorted LLRR order is generated by $[R(\hat{z},-\frac{\pi}{2})T_{a}]^{2}T$ and $R(\hat{y},\pi)T_{a}I$, hence only invariances in Eq. (159,170) are required, leading to
$x_{1}=-x_{2}=x_{3}=-x_{4}$
$y_{1}=y_{2}=y_{3}=y_{4}$
$z_{1}=z_{2}=-z_{3}=-z_{4}.$ (173)
This is the pattern given in Eq. (74).
B.5 FM I order in Kitaev-Heisenberg-Gamma chain
In the Kitaev-Heisenberg-Gamma chain, the unbroken symmetry group $H_{2}^{\text{(I)}}$ for the FM I order is generated by $T_{a}I$ and $R(\hat{n}_{1},\pi)T_{a}$ as discussed in Eq. (86). The spin expectation values transform under these two operations according to
$T_{a}I$ ::: $\left(\begin{array}[]{c}x_{i}\\ y_{i}\\ z_{i}\end{array}\right)\rightarrow\left(\begin{array}[]{c}x_{1-i}\\ y_{1-i}\\ z_{1-i}\end{array}\right),$ (180)
$R(\hat{n}_{1},\pi)T_{a}$ ::: $\left(\begin{array}[]{c}x_{i}\\ y_{i}\\ z_{i}\end{array}\right)\rightarrow\left(\begin{array}[]{c}-y_{i+1}\\ -x_{i+1}\\ -z_{i+1}\end{array}\right).$ (187)
By considering a two-site unit cell, the invariance under $T_{a}I$ requires
$x_{1}=x_{2},\leavevmode\nobreak\ y_{1}=y_{2},\leavevmode\nobreak\ z_{1}=z_{2},$ (188)
and the invariance under $R(\hat{n}_{1},\pi)T_{a}$ requires
$x_{1}=-y_{2},\leavevmode\nobreak\ y_{1}=-x_{2},\leavevmode\nobreak\ z_{1}=-z_{2}.$ (189)
Combining Eq. (188) and Eq. (189), we obtain the pattern of the magnetic ordering in Eq. (87).
B.6 FM II order in Kitaev-Heisenberg-Gamma chain
In the Kitaev-Heisenberg-Gamma chain, the unbroken symmetry group $H_{2}^{\text{(II)}}$ for the FM I order is generated by $T_{a}I$ and $TR(\hat{n}_{1},\pi)T_{a}$ as discussed in Eq. (91). The spin expectation values transform under $TR(\hat{n}_{1},\pi)T_{a}$ according to
$TR(\hat{n}_{1},\pi)T_{a}$ ::: $\left(\begin{array}[]{c}x_{i}\\ y_{i}\\ z_{i}\end{array}\right)\rightarrow\left(\begin{array}[]{c}y_{i+1}\\ x_{i+1}\\ z_{i+1}\end{array}\right).$ (196)
The invariance under $TR(\hat{n}_{1},\pi)T_{a}$ requires
$x_{1}=y_{2},\leavevmode\nobreak\ y_{1}=x_{2},\leavevmode\nobreak\ z_{1}=z_{2}.$ (197)
Combining Eq. (188) with Eq. (197), we obtain the pattern of the magnetic ordering in Eq. (92).